### v3

# Query Rewriting via LLMs

Abstract
When complex SQL queries suffer slow executions despite query optimization, DBAs typically invoke automated query rewriting tools to recommend “lean” equivalents that are conducive to faster execution. The rewritings are usually achieved via transformation rules, but these rules are limited in scope and difficult to update in a production system. Recently, LLM-based techniques have also been suggested, but they are prone to semantic and syntactic errors.
We investigate here how the remarkable cognitive capabilities of LLMs can be leveraged for performant query rewriting while incorporating safeguards and optimizations to ensure correctness and efficiency. Our study shows that these goals can be progressively achieved through incorporation of (a) an ensemble suite of basic prompts, (b) database-sensitive prompts via redundancy removal and selectivity-based rewriting rules, and (c) LLM token probability-guided rewrite paths. Further, a suite of logic-based and statistical tools can be used to check for semantic violations in the rewrites prior to DBA consideration.
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms. The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer. For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison.
Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise query performance.
1 Introduction
The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools. Query optimizers are expected to, in principle, automatically remove such wasteful bloat while constructing efficient execution plans. However, in practice, they are often led astray by complex representations, resulting in poor response times.
SELECT t.Key, sum(t.Rating) AS PostRating,
(SELECT sum(b0.Rating)
FROM (SELECT p0.PostId, p0.BlogId, p0.Content,
p0.CreatedDate, p0.Rating, p0.Title,
b1.BlogId AS BlogId0, b1.Rating AS Rating0,
b1.Url, p0.day AS Key
FROM Posts AS p0 INNER JOIN Blogs AS b1
ON p0.BlogId = b1.BlogId
WHERE b1.Rating > 5) AS t0
INNER JOIN Blogs AS b0 ON t0.BlogId = b0.BlogId
WHERE t.Key = t0.Key ) AS BlogRating
FROM (SELECT p.Rating, p.day AS Key
FROM Posts AS p INNER JOIN Blogs AS b
ON p.BlogId = b.BlogId
WHERE b.Rating > 5) AS t
GROUP BY t.Key;
Figure 1: Complex SQL Representation
SELECT p.day AS Key, SUM(p.Rating) AS PostRating,
SUM(b.Rating) AS BlogRating
FROM Posts AS p INNER JOIN Blogs AS b
ON p.BlogId = b.BlogId
WHERE b.Rating > 5
GROUP BY p.day;
Figure 2: Lean Equivalent Query
As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs. When this seemingly complex query is given to the current PostgreSQL engine (v16) [2], its execution plan essentially mimics the hierarchical structure of the query. This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete. However, the query can be equivalently rewritten (assuming NOT NULL column constraints and key-joins) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools in the hope that they may recommend performant alternatives. For this context, a viable S2S query transformer should satisfy the following criteria: (1) A wide range of queries should be covered in its scope; (2) The transformed query should be semantically equivalent to the original; (3) The rewrite should ideally improve performance, but at least not cause regression; and (4) The transformation overheads must be practical for deployment.
Prior Work
A range of innovative Rule-based (e.g. Learned Rewrite [48]) and Model-based (e.g. Gen-Rewrite [28]) approaches have been proposed for S2S transformations. A recent vision paper [49] also advocates the use of LLMs for such query rewriting. The state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting. However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space (i.e optimization on the nodes of the execution plan tree) rather than directly in query space (i.e transforming the query structure itself). In this paper, we address these lacuna and amplify the effectiveness of query rewriting. In particular, our query space-based transformations provide greater scope for candidate rewrites since only intent, and not implementation, is expressed, and this multiplicity of rewrites in turn results in a broader coverage of the underlying plan space.
Figure 3: High-level architecture of LITHE
The LITHE Rewriter
We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention. As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance. Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$. Note that having the DBA in the loop is a common practice in commercial query advisory systems [14].
Our design of LITHE is based on a calibrated investigation of the suite of techniques described below:
Basic Prompts. We begin with a suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs. Interestingly, we find that more information is not necessarily better wrt rewriting quality, and that the best prompt granularity is query-specific. Moreover, this basic prompt ensemble was found to itself deliver performance similar to the SOTA techniques.
Database-sensitive Prompts. To help the model adapt to different query patterns and structures, we next introduce rules in the prompts. Our rules are invoked directly in query space, providing the LLM with the latitude to generalize the rule usage to a wide range of queries. This is in contrast to the hardwired and narrow rule application mechanisms (e.g. Calcite [10] rules) typically used in existing rewrite systems, which operate in plan space.
In particular, we work with two classes of rules – 1. Redundancy Removal Rules: These rules eliminate repeated and redundant computations of the same output; and 2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
Token Probability Driven Rewrites. The above rewrite options are restricted to the standard prompt interface. We also leverage the rich telemetry provided by LLMs – particularly, the token probabilities output at each step in the prediction sequence. Whenever the LLM lacks high confidence in the next token, we follow multiple alternative paths in the decision process. To ensure practical overheads on this enumerative approach, a Monte Carlo tree search (MCTS) technique is incorporated in our implementation. MCTS has previously been used to improve ordering of query rewrite rules [48] in the plan space, but our usage is new – it is geared to guide the LLM in exploring rewrites in the query space.
Semantic Equivalence Labels. After generating a rewrite, it is imperative to check for semantic equivalence to the original query. Unfortunately, the general problem of proving query equivalence is NP-complete [7]. We therefore take a twofold approach to check for query equivalence. First, the queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database. While these checks do not suffer from false negatives, they may (rarely) incur false positives. These tests can thus quickly weed out obviously non-equivalent rewrites.
A final check is then performed using logic-based equivalence-proving tools that provably work for a restricted classes of queries. Specifically, LITHE uses the recently developed QED tool [40] and SQLSolver [17] to verify rewrite correctness for queries within their scope. Whereas, for out-of-scope queries, result equivalence on the original database itself is used to test the equivalence.
After carrying out the above exercise, LITHE provides a verification label of semantic equivalence – provable if verified by a logic-based tool, or statistical otherwise. Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
Regression Identification. Finally, it is well-known that in practice there are often material discrepancies between optimizer cost predictions and actual execution times [29]. So it is possible that a promising rewrite as per the optimizer may actually regress at run-time as compared to the original formulation. To help guard against such possibilities, we have implemented heuristic mechanisms, described in Section 6.3, to identify “brittle” rewrites and discard them in favor of the original.
Experimental Evaluation
Our first set of experiments to evaluate LITHE’s performance is carried out on the industry standard TPC-DS benchmark [13], hosted on the PostgreSQL platform with GPT-4o used as the LLM. The evaluation focuses on slow queries taking more than a threshold time to complete. We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [48], LLM-R2 [27], GenRewrite [28], as well as a baseline LLM prompt [28]). The primary metrics are (a) reductions in optimizer-estimated costs, (b) run-time speedups, and (c) rewriting overheads. For LLM-based techniques, the number of tokens used is also monitored since the financial charges for LLM usage are typically dependent on this number. To understand the independent utility of the various components of LITHE, a systematic ablation study is carried out.
Generalizability.
In our second stage of experiments, we evaluate generalizability of the above outcomes in a variety of new scenarios, including (a) commercial database engines, (b) unseen database schemas, and (c) alternative LLM platforms.
Results. Our experiments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract costs. In particular, for TPC-DS, LITHE constructed “highly productive” ($>1.5x$ estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number. Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
We also evaluated whether the above cost reductions translated into real execution speedups. Here, we find that LITHE is indeed often substantively faster at run-time as well. Specifically, the geometric mean of the runtime speedups for slow queries was as high as 13.2 over the native optimizer, whereas SOTA delivered 4.9 in comparison. Finally, our regression identification techniques were successful in guarding against “brittle” rewrites that are promising cost-wise, but eventually poor at run-time.
Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise application performance.
Contributions
In summary, our study makes the following contributions:
  1. 1.
    Assesses LLM suitability for S2S transformation.
  2. 2.
    Transforms directly in query space instead of plan space intermediates, leading to performant rewrites.
  3. 3.
    Incorporates database-sensitive rules in LLM prompts, covering both schematic and statistical dimensions.
  4. 4.
    Leverages LLM token probabilities to guide navigation of the rewrite search space and minimize LLM errors.
  5. 5.
    Evaluates rewrite quality over a broad range of database environments, demonstrating substantial benefits over both SOTA and the native optimizer.
  6. 6.
    Identifies learnings that could help guide research directions for industrial-strength query rewriting.
2 LITHE Overview
The LITHE architecture, illustrated in Figure 3, consists of the following five-stage pipeline:
1. Prompt-based rewriting. This component consists of two modules: LLM Prompting and Syntax Verification. The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite. The LLM output’s syntax is then verified with the database engine’s parser– if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively (within a threshold) until a valid SQL query is obtained.
This process is repeated for each of the different prompts described in Sections 3 and 4, and the resulting syntactically-valid rewrites are fed to the query costing component. We evaluate each prompt within a fresh LLM context, thereby making the prompt processing to be order-independent.
2. Query Costing. The costs of candidate rewrites are evaluated via the database engine’s optimizer. Rewrites whose costs are greater than that of the original query are immediately discarded. Whereas, the potentially beneficial rewrites (if any) are checked for semantic equivalence to the original query.
3. Fast Semantic Equivalence. Statistical (result-based) techniques are employed to quickly and cheaply assess the semantic equivalence of a recommended rewrite. If the rewrite is deemed valid by this module, it is returned along with the prompt that generated it; otherwise an invalid label is returned.
4. Token probability-driven rewrite. The prompt producing the most beneficial (and valid) rewrite is used as input to a Monte-Carlo tree search (MCTS)-based procedure to further refine the rewrite quality. Specifically, multiple exploration paths are followed whenever the LLM lacks high confidence in the predicted token (details in Section 5). The query costing and semantic equivalence modules are also used internally within this procedure.
5. Final Checks and Output. The least expensive valid rewrite (as identified by the MCTS module) is tested using a suite of provable (logic-based) techniques to generate a final equivalence label (provable or statistical), and its cost benefit over the original query is computed. Further, the execution “brittleness” is assessed using the robustness heuristics. If no valid rewrite is identified, or if the rewrite is expected to be a regression, the original query itself is returned to the DBA. Whereas, if a beneficial rewrite is recommended, an LLM-generated reasoning for the expected performance improvement is also extracted.
2.1 Performance Framework
We consider a query that takes more than T seconds to complete on the native database engine as a “slow query”, potentially triggering intervention by the DBA. Based on common industry perception (e.g. [5]), T is set at 10 seconds in our study. For this context, we define a Cost Productive Rewrite (CPR) as a rewrite that improves a slow query’s performance by at least 1.5 times wrt the optimizer-estimated cost – this aggressive choice of threshold is so that: (a) there is enough headroom in the optimizer prediction that a runtime regression is unlikely; and (b) the benefits of the rewrite substantively outweigh the transformation overheads.
The overall benefit provided by a rewriting tool is quantified by the number of CPR obtained on the slow query workload. Additionally, we also measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites. Finally, to assess the actual run-time benefits, we evaluate TSGM, the geometric mean (GM) of the response-time speedups obtained by these rewrites.
2.2 Query Micro-benchmark
To motivate the progression of strategies incorporated in LITHE, we create an initial micro-benchmark comprising 10 diverse TPC-DS queries for which we were able to hand-craft high-quality CPRs. These queries are processed on GPT-4o, the popular OpenAI LLM, and the rewrites are evaluated on the PostgreSQL v16 database engine. The human rewrites deliver a CSGM of 11.84, serving as an aspirational target to attain computationally. Later, in Section 7, we extend the evaluation to complete benchmarks.
Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sec 345). The TSGM performance is subsequently discussed in Section 7.
3 Basic Prompts
In this section, we explore the simplest interface to LLMs, namely prompting, for query rewriting. We evaluate four basic prompts, enumerated in Figure 4, which cover a progressive range of instructional detail and test the effectiveness of the LLM’s base knowledge.
Figure 4: Templates used for Basic Prompts
Prompt 1: This is the baseline prompt used in [28], which simply asks the LLM to rewrite a given query to improve performance.
Prompt 2: Explicit instructions are included to maintain semantic and functional equivalence while rewriting.
Prompt 3: Verbose instructions are given to rewrite the query, providing step-by-step guidance to the LLM to think rationally. It is first asked to pick out potential inefficiencies in the input query, and then tasked to identify approaches to address these inefficiencies. Finally, it is instructed to apply the identified solution. Essentially, the prompt tries to make the LLM reason akin to human experts.
Prompt 4: The sequence of instructions in Prompt 3 is split into sub-prompts, and provided to the LLM in an iterative manner instead of all at once. The idea is to break down the complex instructions given in Prompt 3 into digestible steps that help the LLM focus on individual tasks.
Performance
Table 1 shows the performance of the four prompt templates on the micro-benchmark. We find that less than half the rewrites are productive with individual prompts. However, a drill-down shows that the best prompt in the ensemble is query-specific – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them. This ensemble approach raises the CPRs to 6 (Row 5 in Table 1) – however, there remain four queries that are not productively rewritten by these prompts.
Table 1: Performance of Basic Prompts.
Prompt # CPR CSGM
Prompt 1 3 1.99
Prompt 2 2 1.85
Prompt 3 4 2.83
Prompt 4 4 3.00
Prompt Ensemble 6 3.23
SOTA Ensemble 3 2.49
The CSGM, shown in the last column of Table 1, is at most 3 for the individual prompts, while the ensemble reaches 3.23. But these speedups, although productive, are all lower than those delivered by the human rewrites.
Finally, an ensemble of SOTA techniques (described in Section 7) was also processed on the same platform. They delivered 3 CPRs with a CSGM of 2.49 (last row in Table 1), indicating the wide gap between the current reality and what is humanly possible.
4 Database-Sensitive Prompts
As discussed above, basic prompting needs to be improved on two fronts: (1) Ensuring productive rewrites where feasible; and (2) Maximizing the impact of these productive rewrites. To address these issues, we incorporate database domain knowledge. Specifically, we design a one shot-based prompting template, augmented with a set of database-aware rewrite rules. The rules are based on common practices followed by DBAs that are widely applicable, and augmented with precise instructions and useful examples to help guide the LLM in the rewriting process.
As a proof of concept, we explore two categories of rewrites here: (a) Rules that eliminate redundancy in the input queries; and (b) Predicate selectivity-based rules that implicitly guide, via query space reformulations, the query optimizer towards efficient query execution plans. Of course, this basic set of rules can be expanded further, but as shown by our experiments, even this minimal set is capable of delivering substantive improvements over a broad set of database environments.
4.1 Redundancy Removal
There are different types of redundancy that can occur in a SQL query – repeated computations, superfluous filter predicates, unnecessary joins, etc. Rules R1 through R4 in Table 2 are designed to tackle such redundancies. The relevant schematic information (e.g. table names, column names, constraints) required by these rules is also provided in the prompt.
Table 2: Rules for Database-sensitive prompts.
Redundancy Removal Rules
R1 Use CTEs (Common Table Expressions) to avoid repeated computation of a given expression.
R2 When multiple subqueries use the same base table, rewrite to scan the base table only once.
R3 Remove redundant conjunctive filter predicates.
R4 Remove redundant key (PK-FK) joins.
Statistics-based Rules
R5 Choose EXIST or IN from subquery selectivity (high/low).
R6 Pre-filter tables involved in self-joins and with low selectivities on their filter and/or join predicates. Remove any redundant filters from the main query. Do not create explicit join statements.
Figure 5: Templates for Database-sensitive Prompts
The template for such rule-based prompts is shown in Figure 5(a) and includes an example to demonstrate the rule application to the LLM – the specific examples used with our rules are available in the Section E. Note that this prompt template allows for only a single rule to be present in the prompt. This was a conscious design choice because LLMs are often overwhelmed by excessive information given in monolithic form. Therefore, we apply each rule using a separate prompt, finally returning the rewrite providing the best performance improvement.
Performance
The performance improvement achieved on the micro-benchmark by an ensemble that adds the redundancy-removing prompts to the basic set (Section 3) is shown in Table 3. We observe that the CPR increases to 7, and CSGM grows to 6.85.
Table 3: Performance with Redundancy Removal Rules.
Prompt # CPR CSGM
[…]
A natural question here would be whether, while retaining the one-rule-per-prompt design, the rules could be progressively applied with the output of one prompt provided as input to the next, and so on. This approach would benefit queries with multiple types of redundancies. However, it also introduces significant computational overheads due to the vast number of possible rule application orders, making it expensive to explore. So, for simplicity, we have chosen to process them individually rather than cumulatively.
Figure 6: Example Queries illustrating Rule 5
4.2 Selectivity-based Guidance
We now turn our attention to rules whose applicability to a query is conditional on the specific database environment, specifically its statistical aspects. For example, consider the alternative rewrites shown in Figure 6 using the EXIST and IN clauses (highlighted in red), respectively – here the appropriate choice is dictated by the selectivity of the inner subquery – EXISTS for high selectivity values and IN for low values. Rule R5 basically encodes this argument as a rule in Table 2. Similarly rule R6, which pre-filters tables that are involved in self-joins and have low selectivity filter and/or join predicates, is also used to guide a rewrite. Note that a specific instruction to not create explicit joins had to be added in rule R6. This is because in the presence of CTEs, the LLM is prone to schematic confusion regarding which attribute belongs to which table, leading it to construct invalid joins.
The input prompt for these rules, as shown in Figure 5(b), is modified to include the following:
  1. 1.
    Estimated selectivities of columns appearing in the WHERE and JOIN clauses – these values are obtained via calls to the cardinality estimation modules of the query optimizer.
  2. 2.
    Clause rewrite rules and instructions based on statistics.
  3. 3.
    Examples relevant to the chosen rewrite rules.
Performance
The performance improvements following addition of selectivity-guided prompts are shown in Table 4. We observe that CPRs are now obtained for 9 of the ten micro-benchmark queries. Moreover, the resulting CSGM increases to 10.57, quite close to the human target of 11.84.
Table 4: Performance of Metadata-infused Prompts.
Prompt # CPR CSGM
[…]
We note in closing that rules R1 through R6 not only add queries to the productive category, but also deliver greater improvement for those already deemed to be productive via the standard prompts of Section 3. Further, to minimize selection overheads, a classifier could be designed to choose the appropriate rule – this option is examined in Section B.
5 Token Probability Driven Rewrite
A key challenge with LLMs is “hallucinations” – the generation of responses that range from being mildly incorrect to completely made up. This is often due to the output tokens having low confidence, which in turn results in “confusing” the LLM and generating suboptimal outputs [20]. In order to have a robust approach for such cases, we take inspiration from the code generation literature [46]. Specifically, we propose a Monte Carlo Tree Search (MCTS) based decoding approach to search for a sequence of LLM-generated tokens that results in both a valid query rewrite as well as performance improvements. As mentioned earlier, MCTS has been previously used for ordering rule applications in query rewrites [48], whereas our goal is to guide the LLM in exploring query rewrites.
This approach models the problem of query rewriting as a decision tree denoting a Markov Decision Process (MDP) [33]. The root node of the tree corresponds to the initial prompt. An edge from a parent node to a child represents a possible token generated by the LLM and is associated with a value denoting the probability of generating this token given the path taken thus far. Here, each edge can be considered as an action of the MDP. A node is considered terminal if the incoming edge corresponds to the “;” token, signalling end of the textual query.
The state of a node $n$ is represented by the partial rewrite created by following the path from the root to $n$ – it is obtained by concatenating the tokens on this path. The root’s state is an empty rewrite, and each terminal node’s state is a complete rewritten query.
Given the hundreds of thousands of tokens in LLM vocabularies, it may be very expensive (both financially and computationally) to construct the entire tree. It is therefore essential to significantly reduce the token search space while exploring the tree for valid rewrites. This is precisely the purpose of MCTS which applies an Upper Confidence Bound (UCB) heuristic [35] to identify the best paths in a tree without computing the entire tree.
Algorithm 1 Token-augmented Rewrite
1:$root$          # Start State
2:$k$              # Maximum number of child node expansions
3:$\theta$              # Probability threshold for node expansion
4:$iter_{max}$     # Maximum number of iterations
5:$Potential$, $visits$, $V$leftarrow empty Map
6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
7:    $visits[root]$leftarrow $visits[root]+1$
8:    $n_{cur}$leftarrow $root$
9:    # Stage 1: Selection
10:    while $len(n_{cur}.children)>0$ do
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
13:    # Stage 2: Expansion
14:    $expand\leftarrow$ True
15:    while $expand$ and ‘;’ $\notin n_{cur}.state$ do
16:        $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
17:        if $P_{next}[0]\leq\theta$ then
18:           for $token\in tokens_{next}$ do
19:               $n_{new}\leftarrow$ new $Node$ with State $n_{cur}.state\cdot token$
20:               Append $n_{new}$ to $n_{cur}.children$
21:           $expand\leftarrow$ False
22:        else
23:           $n_{cur}.state\leftarrow n_{cur}.state\cdot tokens_{next}[0]$
24:    # Stage 3: Simulation - Expand from $n_{cur}$ to full rewrite
25:    $query\leftarrow$ GreedyExpand($n_{cur}$)
26:    $v\leftarrow$ ComputePotential($query$)
27:    $Potential[query]\leftarrow v$
28:    # Stage 4: Backpropagation
29:    while $n_{cur}\neq\text{Null}$ do
30:        $V[n_{cur}]\leftarrow\max(V[n_{cur}],v)$
31:        $n_{cur}\leftarrow\text{Parent}(n_{cur})$
32:# Return valid rewrite with maximum $Potential>1$
33:if $\exists q\in Potential\mid Potential[q]>1$ then
34:    return $q$ having maximum value of $Potential[q]$
35:else
36:    return the original query
5.1 MCTS Search Process
The pseudocode of the search procedure is shown in Algorithm 1. It consists of four stages – Selection, Expansion, Simulation and Back Propagation – that are repeated across $iter_{max}$ iterations:
1. Selection: The first stage is responsible for identifying the most suitable node of the decision tree that is yet to be expanded (i.e., the tokens corresponding to this node have not yet been processed by the search procedure). Starting from the root, this process greedily selects successive edges (actions) till an unprocessed non-terminal node is reached (Lines 6–8 in Algorithm 1). Actions are picked using a UCB that balances exploration and exploitation. The goal is to pick those actions that have either (1) a higher potential to produce correct and faster rewrites (exploitation); or (2) been selected fewer times in the past (exploration).
Specifically, given a node $n$ and a set of possible actions $a\in A$, the next node in this traversal is chosen as:
$n_{\text{next}}=\operatorname*{arg\,max}_{a\in A}UCB(n,a)$ (1)
where $UCB$ is a heuristic sourced from [35], modified to reflect our formulation where values are associated with nodes rather than edges in the tree. It is defined as follows:
$UCB(n,a)=V(n^{\prime})+\beta(n)*P_{LLM}(a|n.state)*\frac{\sqrt{log(visits[n])}}{1+visits[n^{\prime}]}$ (2)
Here, $n^{\prime}$ is the node reached from $n$ by taking action $a$, and the first component $V(n^{\prime})$ represents the exploitation potential of $n^{\prime}$ to produce correct and faster queries (this notion is formalized below in Stage 3). The second component in the equation represents exploration – it is higher for those child nodes of $n$ that are visited less often. Here, $P_{LLM}$ represents the next token probability, and $visits[n]$ is the number of times $n$ has been visited during the search process. $\beta$ is a function that controls the balance between exploration and exploitation. It depends on two hyperparameters $c_{base}$ and $c$ – a higher value of $c_{base}$ makes the algorithm favor exploitation , whereas a higher value of c increases the incentive to explore. $\beta$ is defined as:
$\beta(n)=log(\frac{visits[n]+c_{base}+1}{c_{base}})+c$ (3)
2. Expansion: The second stage is used to expand the unprocessed node $n_{cur}$ chosen by the Selection stage. Specifically, it retrieves from the LLM the top $k$ probable next tokens from $n_{cur}$’s state (Line 12), and expands the decision tree by adding $k$ new child nodes corresponding to these tokens. To make the expansion tractable, multiple child nodes are added only if the probability of the highest token falls below a threshold $\theta$ (Line 13). In other words, when the highest token probability is below $\theta$, it means that the LLM itself is unsure of what the next token should be and therefore it is worth exploring additional options. On the other hand, if the highest token probability is greater than $\theta$, then the tokens are generated in a greedy fashion from the current node until a point where the LLM is again unsure of the next token, or it reaches a terminal node (i.e., completes a query rewrite).
3. Simulation: Here, we determine the potential value, $V(n_{cur})$, to be assigned to $n_{cur}$. This node is expanded greedily based on the highest-probability tokens until a terminal node is reached (Line 21). Then, the complete rewritten query represented by the terminal node’s state is used to compute the potential (Lines 22–23). For a valid rewrite, $V(n_{cur})$ is equal to the speedup it provides w.r.t the original query. However, if invalid (i.e. syntactically or semantically incorrect), $V(n_{cur})$ is assigned a zero value. After every simulation, the complete rewritten query obtained after the greedy expansion of $n_{cur}$ is cached along with $V(n_{cur})$ in a map, $Potential$.
4. Back Propagation: The $V$ value of the simulation for $n_{cur}$ is back-propagated to all its ancestor nodes, and their values are updated iff the new value is higher than the existing value.
Rewritten Query. At the end of all iterations, $q\in Potential$ with highest value $Potential[q]$ that is greater than 1 is returned as the rewritten query (Lines 29–30). In case no such rewrite exists, implying that all the valid rewrites are slower than the original query, the original query itself is returned (Line 32).
5.2 Input Prompt to MCTS
The root state in Algorithm 1 corresponds to the state just after the prompt is fed to the LLM. One way to use this algorithm is to execute it for all the various prompts discussed in the previous sections, and choose the rewrite that provides the best performance. This, however, is expensive both from the aspect of query rewrite time, as well as the number of LLM tokens used. To minimize these costs, given a user query, the LITHE workflow first selects the prompt yielding the most effective rewrite from among the techniques of Sections 3 and 4. It then employs this prompt to initiate the MCTS-based rewrite. In case no prompt provides a lower-cost rewrite, baseline Prompt 1 of Section 3 is used as the fallback option.
5.3 Performance
LITHE’s performance with MCTS-based rewrites fully matches the human target, yielding 10 CPRs and the maximum CSGM of 11.84. While these gains over just prompting may seem marginal, we wish make two observations: (1) MCTS extracts maximum improvement whererever possible, and (2) As shown in Section7.5.2, its impact becomes significant for smaller LLMs like LLaMA.
6 Implementation Choices
In this section, we briefly discuss the design choices made in our implementation of LITHE.
6.1 LITHE Parameter Settings
The “temperature” parameter of GPT-4o, which ranges over [0,1], controls randomness of the model’s response. While higher values suit creative tasks that require diverse and exploratory outputs, we seek focused, deterministic answers. Thus, we set the temperature to 0, forcing the model to greedily sample the next token.
The hyperparameters used by LITHE for MCTS are as follows: The maximum number of iterations $iter_{max}$ is set to 8, expansion threshold $\theta$ is 0.7, and number of expansions $k$ is 2. The values of $c_{base}$ and $c$ were set to 10 and 4, respectively. These settings were determined after an empirical evaluation of the various trade-offs, providing a robust balance between efficiency and quality.
Finally, we try a maximum of 5 times to fix, via prompt corrections, any rewrite that exhibits syntax errors (Section 2).
6.2 Query Equivalence Testing
We use a multi-stage approach, described below, to help the DBA test semantic equivalence between the original query and a recommended rewrite. Note that, as mentioned earlier, having the DBA in the loop is a common practice in commercial query advisory systems[14].
1. Result Equivalence via Sampling. We use a sampling-based approach to quickly test equivalence in the rewrite generation stages of the pipeline. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not sufficient. That is, false positives may be present because the sampled database may not cover all the predicates featured in the query. To minimize this likelihood, we use a combination of (1) correlated sampling [45] for maintaining join integrity in the sample, (2) adjusting constants in the filter predicates to produce populated results, and (3) injecting query-specific synthetic tuples in the sample, similar to the XData mutant-killing tool [34], to cover predicate boundary conditions – the complete details are in Section C.
2. Logic-based Equivalence. Although verifying the equivalence of a general pair of SQL queries is NP-complete [7], a variety of logic-based tools (e.g. Cosette[12], SQL-Solver [17], VeriEQL [21], QED [40]) are available for proving equivalence over restricted classes of queries. Once the least-cost rewrite is identified, LITHE uses the recently proposed QED [40] and SQLSolver [17] since these two approaches were found to cover a larger set of queries compared to the alternatives. The advantage of such a logic-based approach is that it is definitive in outcome. Note that this rewrite has already passed the sampling-based tests described above.
3. Result Equivalence on the Entire Database. If the logic-based test is inconclusive, result equivalence is evaluated on the entire database itself. The DBA may choose to prematurely terminate this test in case the checking time is found to be excessive.
6.3 Regression Identification
We use heuristics to predict whether the actual execution time of a given promising rewrite may turn out to be slower than the original query. The first heuristic is based on the SEER robust plan identification algorithm [19], which is predicated on the database engine providing a “force-plan” feature (e.g. [1]). Let $P^{U}$ and $P^{T}$ be the plans corresponding to the original query, $Q^{U}$, and the recommended rewrite, $Q^{T}$, respectively. We construct parametrized versions of these queries, where the constants in the filter predicates are replaced by variables. Then, by assigning appropriate values to these parameters, we construct queries that are located at the corners of the selectivity space. The plans $P^{U}$ and $P^{T}$ are forced at each of these corner locations, and if the rewrite’s cost is lower than the original at all of them, the rewrite is deemed to be robust.
At this time, plan forcing is not available in all database engines. For such limited environments, we fall back to an alternative heuristic – we compare the query runtimes obtained on the series of executions carried out on sampled databases during semantic equivalence testing. If the rewrite’s runtimes are lower in all of these intermediate executions, then it is estimated to be robust.
7 Experimental Evaluation
In this section, we report on LITHE’s performance profile. We first describe the experimental setup, including comparative baselines, query suites and evaluation platforms. Then we present the speedup results for both aggregate benchmark and individual queries, followed by characterization of the rewrite overheads in computational and financial terms. We finally discuss the impact of alternative platforms wrt database engine, database schema and LLMs.
Rewrite Baselines. We compare LITHE with a collection of contemporary rewrite techniques, collectively referred to as SOTA – the details of these techniques are provided in Section 9. Specifically, the SOTA collection consists of the following approaches:
  1. 1.
    Baseline LLM prompt [28]: This is Prompt 1 from Section 3.
  2. 2.
    Learned Rewrite [48], a purely rule-based rewriter.
  3. 3.
    LLM-R2 [27], an LLM-guided rule-based rewriter.
  4. 4.
    GenRewrite [28], a purely LLM-based rewriter.
Given an input query, each of the SOTA approaches is independently invoked to perform a rewrite, and the rewrite with the best performance is used as the baseline for comparison. Note that these approaches may occasionally generate rewrites that are expected by the optimizer’s costing module to regress the performance. For safety, we immediately discard such rewrites, similar to LITHE.
Query Set. Our evaluation in this paper primarily focuses on complex analytical queries from the standard TPC-DS decision-support benchmark [13], which models a retail supplier environment. The benchmark is used at its default size of 100 GB. As mentioned earlier, we focus on “slow queries” that take over 10 seconds to execute in their original form, operating in a cold-cache environment.
LITHE has also been evaluated on other benchmarks, including DSB [15], ARCHER [47], JOB [23] and StackOverflow [30]. The overall performance characteristics were found to be similar to those presented here – due to space limitations, we defer their details to Section A.
Testbed. The majority of our experiments were carried out on the following data processing platform: Sandbox server with Intel(R) Xeon(R) CPU E5-1660 v4 @ 3.20GHz x 16, 32 GB RAM, and 12TB HDD, running Ubuntu 22.04 LTS; PostgreSQL v16 database engine; and GPT-4o LLM for both LITHE and SOTA. Variations on this platform are considered in Sections 7.4 and  7.5.
Metrics. For each rewrite technique, we identified the number of queries for which a CPR (cost productive rewrite with $>$ 1.5 speedup) could be constructed. Subsequently, we computed the CSGM (Cost Speedup Geometric Mean) and TSGM (Time Speedup Geometric Mean) performance obtained by each technique over the set of all CPRs (i.e. CPRs arising from either LITHE or SOTA).
From the investment perspective, we measured the average rewrite time per query, and additionally for the LLM-based techniques, the number of tokens used in the rewrite process.
7.1 Rewrite Quality (Cost and Time)
7.1.1 Estimated Cost
LITHE produces a rewrite with a positive cost speedup ($>$ 1x) for 46 of the 88 TPC-DS queries deemed to be slow by our threshold. Of these 46, there were 26 CPRs resulting in a highly productive CSGM of 11.5. On the other hand, SOTA delivers only 13 CPRs (out of 42 positive rewrites) with a CSGM of 6.1. All but one of the SOTA CPRs also feature in the LITHE CPRs, making the total number of CPRs considered to be 27. Of these 27, we were able to formally verify 11 using the logic-based tools, whereas the remaining 16 passed our statistical tests. Furthermore, we also manually verified the correctness of these rewritten queries.
A drill-down into the cost speedup performance at the granularity of individual queries is shown in Figure 8, which compares LITHE (orange bars) and SOTA (blue bars) on each of the 27 CPR queries – note that the cost speedups on the $x$-axis are tabulated on a $\mathbf{log_{10}}$ scale, and the queries are sequenced in decreasing order of LITHE speedup. The vertical dotted line at $1$ represents the normalized baseline cost of the original query with the native optimizer, while the vertical line at $1.5$ is the CPR threshold.
We first observe, gratifyingly, that rewrites are indeed capable of promising dramatic cost speedups – take, for instance, Q41, which improves by a whopping five orders-of-magnitude for both SOTA and LITHE. This improvement in query performance is due to replacing the “WHERE (SELECT COUNT(*) from ...) > 0” clause with “WHERE EXIST (SELECT 1 from ...)” – the latter is a more efficient check for result existence in an inner subquery since it removes the computationally expensive aggregation function.
Figure 7: Plan Cost Speedups
Figure 8: Execution Time Speedups
Second, in most queries, LITHE’s cost speedup either exceeds or matches SOTA. In fact, for several queries (e.g. Q45, Q25, Q4), LITHE produces a highly beneficial CPR but SOTA returns the original query. Conversely, the opposite is true for one query (Q57) where SOTA projects a large speedup but LITHE settles for the original query. And in Q88 and Q95, SOTA performs only marginally better.
At this stage, one might expect that adding more rules to LITHE could bring it on par with SOTA for queries like Q57. However, we deliberately include only broad-brush rules in LITHE to ensure generalizability and efficiency. As the following timing section shows, due to this conservative approach LITHE actually outperforms SOTA on these queries (Q57, Q88, Q95) in terms of runtime.
7.1.2 Execution Time
Thus far, we had considered optimizer-estimated execution costs. We now move to wall-clock runtimes for query executions – Figure 8 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA on a cold-cache environment. We observe, again gratifyingly, that there are indeed several queries where substantial time benefits are achieved by the rewrites, even exceeding order-of-magnitude benefits in some cases – for instance, LITHE improves Q45 by a large factor of 700! Second, in all cases, LITHE outperforms or matches SOTA, including as mentioned above, the queries where SOTA’s optimizer costs were better.
From a modeling perspective, we see that the well-documented gap between optimizer predictions and actual run-times is prevalent in the rewrite space as well. On the one hand, there is Q45 where the projected speedup of 40X increases to 700X at runtime. When run in a warm-cache environment, this speedup increases to a huge $10^{4}$ (see Section A). On the other hand, the $10^{5}$ speedup for Q41 decreases to a still reasonable 200X. But for SOTA, the reductions can be severe – a striking case in point is Q57, where SOTA actually causes regression despite a speedup projection of close to 100. However, the good news is that with LITHE, although the runtime speedups did not always match the projections, we did not encounter any regressions among the CPR rewrites. Overall, LITHE produces more robust rewrites resulting in a high TSGM of 13.2, whereas SOTA provides a TSGM of 4.9.
7.1.3 Reasoning
As a confirmatory exercise, we compared the LLM-generated explanation output by LITHE with our own manual analysis of the query plans generated for the original and rewritten queries. For example, consider Q45 – the provided explanation correctly lists the main reasons for the 700X speedup as “the use of CTEs”, “Reduction in data volume early on” through pre-filtering based on join predicates, etc. Overall, we found that the explanations provided by LITHE matched our manual analysis of the plans, indicating that model-based reasoning is well aligned with human-backed reasoning in these scenarios.
7.2 Ablation Analysis of Rewrite Quality
7.2.1 Components of LITHE
A natural question at this stage is the role of the various techniques in LITHE towards achieving its large performance benefits. This analysis is captured in Table 5 which lists the CPRs for each technique when invoked in isolation as well as the cumulative number of CPRs when the different techniques are combined in the order of listing.
Table 5: CPRs contributed by LITHE components.
# CPR
Isolated Cumulative
Basic Prompts (Section 3) 15 15
Rules R1 — R4 (Section 4.1) 10 18
Rules R5, R6 (Section 4.2) 11 25
MCTS (Section 5) 26 26
We observe that the Basic Prompt ensemble and Redundancy Rules (R1 – R4) contribute to around two-thirds of the CPRs (18/26). When the statistics-infused Rules (R5, R6) are added, this number jumps to 25. Finally, MCTS produces an additional CPR and also improves the cost speedup of a pre-existing CPR.
7.2.2 Database-Sensitive Prompts.
Since MCTS explores the search space at a fine granularity, one could ask whether just the Basic Prompts in conjunction with MCTS would suffice to provide good performance. The motivation is that it would relieve us from using the database-sensitive rules that incur significant computational and financial overheads. When this experiment was conducted, the CSGM dropped precipitously to a paltry 5, a far cry from the 11.5 obtained with the database-sensitive rules. These results highlight the need to reflect database awareness for effective query rewriting, and not rely solely on prior LLM knowledge.
7.3 Rewrite Overheads (Time/Money)
Having established the performance benefits of rewrites, we now turn our attention to their time and financial overheads.
7.3.1 Transformation Time
The average processing time per CPR query is shown in Table 6, and we see that they do take a few minutes. However, this investment may be acceptable in deployment given that the execution benefits typically far outweigh the compilation overheads. For instance, with Q11, the original query took nearly an hour to complete, whereas the LITHE rewrite executed in just 3 minutes. Further, many applications tend to use a set of canned queries which are run thousands of times. Thus, even a large one-time investment can be easily recovered over repeat executions of such queries.
Table 6: Rewrite Overheads of LITHE and SOTA.
Avg. Time (min) Avg. Tokens Avg. Cost (USD)
LITHE 5 18427 0.045
SOTA 1.7 20076 0.050
Notwithstanding the above, we also observe that LITHE is considerably slower than SOTA in producing rewrites. A drill-down shows that the lion’s share of the time is taken by the initial prompt ensemble and the final MCTS module. We explore techniques to improve their efficiency in Section B.
7.3.2 Monetary Outlay
The average number of LLM tokens required by LITHE and SOTA, and their associated financial costs [^1] At the time of writing, GPT-4o costs USD 2.5 per million tokens., are also shown in Table 6. The good news is that the inference charges per query are just a few cents, making rewriting practical from a deployment perspective.
7.4 Commercial DBMS
A legitimate question could be whether the rewrites made amends for the PostgreSQL optimizer but may fail to be useful in highly-engineered database engines. To evaluate this issue, we performed TPC-DS rewrites on a pair of popular commercial DBMS, which we refer to as OptA and OptB.
Table 7: Performance on Commercial Database Engines.
# CPR CSGM TSGM
OptA OptB OptA OptB OptA OptB
LITHE 12 9 3.6 4.1 2.1 1.9
SOTA 3 5 1.5 1.3 1.4 1.2
The performance of LITHE and SOTA on these two systems is shown in Table 7, with LITHE continuing to do better than SOTA. Despite the apparent lack of optimization headroom, LITHE still produces 12 and 9 CPRs resulting in a healthy CSGM of 3.6 and 4.1, respectively. Further, the TSGM provided by these rewrites are a useful 2.1 and 1.9, respectively.
Interestingly, although we did not observe any regressions with PostgreSQL, a few did surface in the commercial systems. Nevertheless, our regression identification mechanisms effectively caught these brittle rewrites. As a case in point, a promising rewrite, as estimated by the optimizer, for Q23, actually takes 37 minutes to complete as compared to 18 minutes for the original query – this doubling slowdown was successfully flagged by the SEER heuristic, and the rewrite was abandoned.
We also conducted a preliminary study of LITHE on an internal real-world benchmark with more realistic data and query characteristics than TPC-DS. Even on this benchmark, LITHE produced 8 CPRs (out of 45 queries tested) with CSGM of 2 and TSGM of 3.3.
The above results suggest LITHE has a useful role to play in industrial environments. From a different perspective, a company building a new database engine could use LITHE to non-invasively overcome the limitations of early versions of its optimizer.
7.5 Dependence on LLM (Training/Model)
7.5.1 Masked Database
An interesting question to ask now is whether the performance benefits seen thus far could be an artifact of GPT-4o having already been trained well on the TPC-DS benchmark, which is prominent in the public domain. To investigate this issue, we created a masked version of the TPC-DS database schema, whereby the table and column names convey no semantic information about their contents. We then constructed rewrites for the CPR queries (after syntactic changes to reflect the new masked schema) on this version. The results are shown in Table 8, and we observe that the performance profiles only marginally decrease for both LITHE and SOTA, testifying to their robust applicability.
Table 8: Rewrite Performance on Masked Database.
Approach # CPR CSGM TSGM
TPC-DS Masked TPC-DS Masked TPC-DS Masked
LITHE 26 24 11.5 10.5 13.2 11.8
SOTA 13 12 6.1 5.3 4.9 4.2
7.5.2 LITHE on LLaMA
In our concluding experiment, we evaluated LITHE’s performance on the LLaMA 3.1 70 billion parameter instruct model, substantially smaller compared to GPT-4o, which may have several hundred billion parameters [6]. To attain practical inference times, the model was loaded using a low 4-bit quantization. Further, to ensure reproducibility and deterministic answers, the do_sample parameter was set to False, which forces the LLM to perform greedy decoding. To make up for the huge reduction in model parameters as compared to GPT-4o, we include up to two example demonstrations for each rule-based prompt.
For this environment, Table 9 shows LITHE’s performance on the TPC-DS benchmark with and without MCTS. Although certainly lower than the corresponding numbers with GPT-4o (Section 7.1), it is encouraging to see that, in absolute terms, significant performance benefits can be obtained for most queries, especially with MCTS support. So, the message is that smaller models can also be fruitfully used in real-world environments.
Table 9: LITHE Rewrite Performance with LLaMA
# CPR CSGM TSGM
LLaMA without MCTS 18 5.6 6.5
LLaMA with MCTS 22 8.5 10.9
8 The Road Ahead
Based on our study, we now present a few observations with implications for the future design and deployment of rewriting tools.
8.1 Rewrite Space Coverage by LLMs
Given the decades-long research on database query optimization, we expected the potential for performance improvement via rewriting to be limited. What came as a surprise was the substantial scope for improvement still available, as showcased by the large CSGM and TSGM values, even on commercial platforms. These results suggest that LLMs explore optimization spaces that are well outside the purview of contemporary database engines. Further, this enhanced space could be augmented, in a two-stage process, with the recent proposals for LLM-based “plan hints” that steer the optimizer in fruitful directions within a plan space [8].
8.2 Rewrite Migration to Optimizer
The above demonstrated the potent exploratory power of LLMs. But from an overheads perspective, such rewrites should ideally be within the optimizer’s native search space rather than recommended from outside. Therefore, it would be a useful exercise to try and distill fresh optimization rules from these instances, leveraging the extensibility features of contemporary optimizers [16] to facilitate their incorporation in existing systems.
Figure 9: Rewritten TPC-DS Q90
On the flip side, there appears to be an “impedance mismatch” against such integration for certain classes of rewrites. For example, consider the TPC-DS Q90 rewrite in Figure 9. The original query individually computed AM (morning) sales and PM (evening) sales, which were then used to compute the AM to PM ratio. The rewrite, however, extracted all relevant rows in one shot and computed the ratio using CASE statements – encoding such transformations as generic rules in the optimizer appears challenging, given the combinatorial ways in which such transformations can occur.
Therefore, a fruitful area of future research could be achieving a middle-ground between the disparate world-views of LLMs and traditional optimizers.
8.3 Agentic LLMs for Query Rewriting
An effective way to extend LITHE is to use an agentic LLM that actively interacts with the database environment. This would allow a “LITHE agent” to leverage a memory store to log one or more of: (1) query execution times; (2) memory usages of slow queries; (3) prior interactions with users; and (4) useful/new rewrite rules learnt over time. Additionally, the agent could be given database engine access so that it can directly request metadata such as query execution plans, database statistics, live query analyzers, etc. By following a structured workflow – planning, acting, observing, and refining – the agent would then iteratively enhance query performance. This adaptive workflow supports continuous learning, so that query rewriting can become more efficient over time.
8.4 Scope of Semantic Equivalence Tools
As seen in the experiments section, logic-based query equivalence testing covers industrial-strength queries only to a limited extent. On the other hand, while it is highly likely that the statistics-verified rewrites are valid, it still requires the DBA to make a final call on the correctness. This limitation restricts the use of LITHE in a fully automated scenario, i.e., as a direct preprocessor to the query engine. Therefore, a key challenge is to improve logic-based coverage.
9 Related Work
Rule-based SQL rewriting. Most of the recent work on SQL query rewriting is rule-based [48, 42, 9, 43, 11, 31]. For instance, WeTune [42] uses a rule generator to enumerate a set (up to a maximum size) of logically valid plans for a given query to create new rewrite rules, and uses an SMT solver to prove the correctness of the generated rules. While this approach can generate a large set of new rewrite rules, it often fails in coming up with transformation rules for complex queries due to the computational overheads of verifying rule correctness. As such, it is unable to rewrite any of the TPC-DS queries [18, 28].
Learned Rewrite [48] uses existing Calcite [10] rules and aims to learn the optimal subset of rules along with the order in which they must be applied. Since the rewrite search space grows exponentially with the number of rules, it uses MCTS scheme to efficiently navigate this space and find the rewritten query with maximum cost reduction. This is in contrast to LITHE where MCTS is used to search over the output token space of an LLM in order to prevent LLM hallucinations. Specifically, our use of MCTS ensures that the LLM correctly follows the prompt instructions.
LLM-R2 [27] is also rule-based but takes a different approach to identify the order for rewrite rule applications: it uses an LLM to find the best Calcite rules and the order in which to apply them to improve the query performance. R-Bot [36] also leverages an LLM to optimize the order of Calcite rules, but employs advanced contemporary techniques such as retrieval-augmented generation (RAG) and step-by-step self-reflection to improve the outcomes.
Query Booster [9] implements human-centered rewriting – it provides an interface to specify rules using an expressive rule language, which it generalizes to create rewrite rules to be applied on the query. There are also rule-based rewrite approaches designed for specific types of rewrites such as optimizing correlated window aggregations [43] and common expression elimination [11].
All of the above approaches operate via the query plan space, which can restrict the kind of rewrites that can be accomplished. Whereas, LITHE uses a small set of general rewrite rules that work directly in the query space.
LLM-based rewriting. GenRewrite [28] is the first LLM-based approach to use the LLM for end-to-end query rewriting. Instead of using predefined rules from Calcite [10], they employ the LLM to create Natural Language Rewrite Rules (NLR2s) to be used as hints, and perform several iterations of prompting to get the rewritten query. They show that LLMs can outperform rule-based approaches due to their ability to understand contexts, and demonstrate a significant improvement in query rewriting compared to prior methods.
A limitation, however, is that LLM-generated rewrite rules often fail to generalize beyond specific query pairs. Even when generalized rules are present, LLMs can struggle to apply rules correctly if not provided with accompanying examples. Finally, it must be noted that LLMs are unaware of the underlying database which restricts their ability to produce efficient metadata-aware rules.
LLMs for Database Modules. LLM technologies have been advocated for a variety of database modules. For instance, they have been extensively used for Text-to-SQL transformations [26, 44, 3, 4, 37, 49]. The main focus of these techniques is to correctly ascertain the information necessary to formulate the SQL query [39, 32, 37]. On the other hand, the goal of S2S rewriting is on improving the performance of an existing SQL query. Thus, unlike Text-to-SQL transformations where the input text is inherently ambiguous, SQL queries are precisely defined, and therefore equivalence to a precise ground-truth has to be provably maintained.
In recent times, LLMs have also been considered for plan-hinting [8], join-order optimization [38], index selection [49], data pipelines [22], data management [25], etc. An LLM-enabled multi-modal query optimizer [41], where data spans text, image, audio and video domains, has also been proposed recently. These approaches can be used in conjunction with the LITHE system since they address orthogonal segments of the query processing pipeline.
10 Conclusions and Future Work
We investigated how the latent power of LLM technologies can be productively materialized in the context of SQL-to-SQL rewriting. Our study progressively infused database domain knowledge, such as redundancy removal rules and schematic+statistical metadata, into the LLM prompts. Further, the output telemetry of LLMs, in the form of token probabilities, was used to signal situations where the LLM lacked confidence, triggering exploration of a larger search space. Finally, a combination of logic-based and statistical tests was employed to verify the equivalence of the rewrites.
An empirical evaluation over common database benchmarks showed that rewriting is a potent mechanism to improve query performance. In fact, even order-of-magnitude speedups were routinely achieved with regard to both abstract costing and execution times. However, our results also showed a significant semantic distance between foundation models and query optimizers, with regard to both scope and precision, which would have to be bridged to fully leverage the latent power of LLMs. Further, our focus here was primarily on prompting-based strategies – in our future research, we plan to investigate how domain-specific fine-tuning could be leveraged to provide GPT-4o-like rewrites on small open models.
Appendix
Appendix A Additional Experiments
A.1 Rewrite Quality (TPC-DS) Contd.
Figure 10 shows the runtime speedups (on a $\mathbf{log_{10}}$ scale) obtained by LITHE and SOTA for the queries run on a warm-cache environment.
Figure 10: Execution Time Speedups in a warm-cache scenario.
As expected, the speedups often significantly increase when compared to the cold-cache scenario. For instance, the speedup for Q45 increases from 700X to over $10^{4}$. Overall, the TSGM of LITHE increases to 17.1 (from 13.2), while that of SOTA increases to 5.7 (from 4.9).
A.2 Rewrite Quality (Other Benchmarks)
We also evaluated LITHE on the following set of benchmarks.
  1. 1.
    DSB [15]: A TPC-DS variant with complex data distributions and additional query templates featuring many-to-many joins and non-equi-joins.
  2. 2.
    ARCHER [47]: A Text-to-SQL benchmark spanning 10 databases with availability of ground-truth SQL queries.
  3. 3.
    JOB [24]: An optimizer stress-test benchmark featuring queries with large and complex join graphs.
  4. 4.
    StackOverflow [30]: A real-world benchmark with query templates modeling questions and answers from experts. A random instance of each template is taken.
Note that the queries in these workloads (with the exception of DSB) are mostly fast running (i.e. completing in less than 10 seconds). Therefore, with the primary goal to test the coverage of LITHE over a diverse variety of workloads (and not the deployment angle), we include all the queries in our analysis. The number of CPRs produced and the cost speedup delivered by LITHE and SOTA over these benchmarks are shown in Table 10.
Table 10: Comparing LITHE with SOTA on different benchmarks.
Benchmark # CPR CSGM
LITHE SOTA Union LITHE SOTA
DSB 9 3 9 7.7 1.7
ARCHER 22 19 22 2.1 1.95
JOB 4 2 4 1.9 1.4
StackOverflow 2 1 2 8.7 7.5
In case of DSB, LITHE produces CPR for 9 queries resulting in a highly productive CSGM of 7.7. Similar to the case with TPC-DS, LITHE performs better than SOTA both with respect to its coverage, as well as the cost speedups.
Turning our attention to the other benchmarks (ARCHER, JOB, StackOverflow), the number of CPR queries is smaller due to the predominance of flat SPJ formulations in these benchmarks, which limits the scope for productive rewriting. Nevertheless, LITHE continues to achieve better CPR coverage, whereas SOTA misses quite a few opportunities. Further, the CSGM of LITHE is visibly better than SOTA.
Appendix B Overheads Reduction for LITHE
Rule Selection Classifier
Reducing the query rewrite time while still obtaining the same performance would be possible if we could directly use the MCTS-driven rewrite with the appropriate prompt. Towards this end, we build a classifier to pick the most appropriate rewrite rule for a given input query. Specifically, the classifier identifies which, if any, of Rules R1–R6 is appropriate for a given query. If none are appropriate, then it falls back to just the set of basic prompts to identify the best prompt to be given as input to the MCTS module. In addition to reducing the rewrite times, using a classifier can also reduce the financial costs of the rewrite.
We design an LLM based classifier to accomplish this as follows: The LLM is given the rewrite rules discussed so far, and additionally, for each rule, an example demonstrating when the rule can be applied and a counter-example demonstrating when the rule cannot be applied. For the database schema and statistics-based rules, the relevant information is also fed to the classifier so that it can make an informed decision. Then, given an input query, the classifier is tasked with selecting the most appropriate rewrite rule.
Table 11: Impact of Classifier (TPC-DS)
Metrics Without Classifier With Classifier
# CPR 26 23
CSGM 11.5 8.5
Avg. Tokens 18427 16003
Avg. Time 5 min 2.4 min
Table 11 compares the performance of LITHE with and without the classifier. The time overheads do visibly go down by about 52 percent, and the tokens by about 13 percent. However, there is a price to be paid – the CPR is reduced to 23 and the CSGM comes down to 8.5. In our future work, we plan to look into whether a better tradeoff could be achieved between quality and overheads.
Pruning in MCTS
A bottleneck in the MCTS-based exploration is the need to greedily expand a node (during the simulation stage) until an entire rewrite is output. In principle, if we could quickly check for semantic and syntactic correctness at intermediate stages, then unproductive paths could be terminated early. We are currently working on the design and implementation of such checks.
Appendix C Query Equivalence Testing via Sampling
We use a sampling-based approach to quickly test equivalence in the rewrite generation stages of the pipeline. The idea here is to execute the queries on several small samples of the database and verify equivalence based on the sample results. However, while this test is a necessary condition for query equivalence, it is not a sufficient condition. That is, there are no false negatives, but there can be false positives. This is because the sampled database may not cover all the predicates present in the query. This can cause two types of problems: First, it is possible for two different queries to return the same result. This can happen when, for example, the entirety of the sampled data satisfies a predicate of one query, while the same predicate is not present in the other. Second, if the underlying sample does not satisfy any of the predicates in either query, then an empty result will be returned by both queries. This again does not imply that the queries are equivalent.
To statistically address the first problem (false positives), we create multiple samples of the database with different seeds, and run the test on all these samples. The goal is to reduce the likelihood of non-equivalent queries returning the same results.
To minimize the occurrence of the second problem (empty results), the following approach is taken:
  1. 1.
    We use correlated sampling [45] to sample the database. This technique leverages the join graph of the schema to produce a sample that maintains join integrity between the tables participating in the query.
  2. 2.
    Given a pair of queries to test for equivalence, we adjust the constants in the filter predicates to reduce the chances of an empty result. We make use of the rows in the sampled data for this purpose. For example, say an equality predicate is present in the query and the associated constant is absent in the sampled database. We then replace the query constant with a value already present in the sample. Similarly, the constants for other comparison operators are adjusted based on the ranges of the corresponding columns in the sampled database.
    Note that these modifications only change the selectivity of the query, but not its semantics.
  3. 3.
    To cover predicate boundary conditions, we inject query-specific synthetic tuples into the sample—similar to the approach used by the XData mutant-killing tool [34]—in order to reduce the likelihood of non-equivalent queries producing the same results.
Appendix D Detailed MCTS Algorithm
Algorithm 2 Token-augmented Rewrite
1:$root$,          # Start State;
2:$k$,              # Maximum number of child expansion
3:$\theta$,              # Probability threshold for node expansion
4:$iter_{max}$,     # Maximum number of iterations
5:$Potential$, $visits$, $V$leftarrow empty Map
6:for $i\leftarrow 1,2,\dots,iter_{max}$ do
7:    $visits[root]$leftarrow $visits[root]$ + 1
8:    $n_{cur}$leftarrow $root$
9:    # Stage 1: Selection
10:    while $len(n_{cur}.children)>0$ do
11:        $n_{cur}$leftarrow $\operatorname*{arg\,max}_{a\in Actions(n_{cur}.children)}$ UCB$(n_{cur},a)$
12:        $visits[n_{cur}]$leftarrow $visits[n_{cur}]+1$
13:    # Stage 2: Expansion
14:    $expand\leftarrow$ True
15:    while $expand$ do
16:        $tokens_{next}$, $P_{next}$leftarrow Model($n_{cur},k$)
17:        if $tokens_{next}[0]=$‘;’ then
18:           $expand\leftarrow$ False
19:        else if $P_{next}\leq\theta$ then
20:           for $token\in tokens_{next}$ do
21:               $state\leftarrow n_{cur}.state\cdot token$
22:               $n_{new}\leftarrow$ new $Node$ with State $state$
23:               Append $n_{new}$ to $n_{cur}.children$
24:           $expand\leftarrow$ False
25:        else
26:           $state\leftarrow n_{cur}.state\cdot tokens_{next}[0]$ First Token
27:           $n_{new}\leftarrow$ new $Node$ with State $state$
28:           Append $n_{new}$ to $n_{cur}.children$
29:           $n_{cur}\leftarrow n_{new}$
30:    # Stage 3: Simulation
31:    $query\leftarrow$ GreedyExpand($n_{cur}$)
32:    # Compute the potential of query (returns 0 for invalid query)
33:    $v\leftarrow$ ComputePotential($query$)
34:    $Potential[query]\leftarrow v$
35:    # Stage 4: Backpropagation
36:    $n\leftarrow n_{cur}$
37:    while $n\neq$ Null do
38:        if $v>V[n]$ then
39:           $V[n]\leftarrow v$
40:        $n\leftarrow\text{Parent}(n)$
41:# Return valid rewrite with maximum $Potential>1$
42:if $\exists q\in Potential\mid Potential[q]>1$ then
43:    return $q$ having maximum value of $Potential[q]$
44:else
45:    return the original query
Appendix E Examples used in Prompts for Rules 1–6
R1: Use CTEs (Common Table Expressions) to avoid repeated computation.
Original Query
SELECT emp.employee_name,
       mgr.manager_name
FROM   employees emp,
       managers mgr
WHERE  emp.manager_id = mgr.manager_id
       AND emp.employee_id IN (SELECT manager_id
                               FROM   (SELECT manager_id,
                                              manager_name
                                       FROM   managers
                                       WHERE  job_id = ’IT_PROG’
                                              AND manager_id > 100))
       AND mgr.manager_name IN (SELECT manager_name
                                FROM   (SELECT manager_id,
                                               manager_name
                                        FROM   managers
                                        WHERE  job_id = ’IT_PROG’
                                               AND manager_id > 100));
Rewritten Query
WITH cte
     AS (SELECT manager_id,
                manager_name
         FROM   managers
         WHERE  job_id = ’IT_PROG’
                AND manager_id > 100)
SELECT emp.employee_name,
       mgr.manager_name
FROM   employees emp,
       managers mgr
WHERE  emp.manager_id = mgr.manager_id
       AND emp.employee_id IN (SELECT manager_id
                               FROM   it_prog_managers)
       AND mgr.manager_name IN (SELECT manager_name
                                FROM   it_prog_managers);
R2: When multiple subqueries use the same base table, rewrite to scan the base table only once.
Original Query
SELECT (SELECT Avg(salary)
        FROM   employees
        WHERE  department = ’Sales’
               AND experience_years BETWEEN 1 AND 5
               AND salary BETWEEN 50000 AND 60000) AS Sales_Avg,
       (SELECT Avg(salary)
        FROM   employees
        WHERE  department = ’HR’
               AND experience_years BETWEEN 5 AND 10
               AND salary BETWEEN 80000 AND 90000) AS HR_Avg;
Rewritten Query
SELECT avg(
       CASE
              WHEN department = ’Sales’ THEN salary) AS sales_avg,
       avg(
       CASE
              WHEN department = ’HR’ THEN salary) AS hr_avg
FROM   employees
WHERE  (
              department = ’Sales’
       AND    experience_years BETWEEN 1 AND    5
       AND    salary BETWEEN 50000 AND    60000)
OR     (
              department = ’HR’
       AND    experience_years BETWEEN 5 AND    10
       AND    salary BETWEEN 80000 AND    90000);
R3: Eliminate overlapping subqueries.
Original Query
SELECT c.*
FROM   customer c
WHERE  c.address_id IN (SELECT a.address_id
                        FROM   address)
       AND c.address_id IN (SELECT a.address_id
                            FROM   address
                            WHERE  a.pin_code = ’560012’);
Rewritten Query
SELECT c.*
FROM   customer c
WHERE  c.address_id IN (SELECT a.address_id
                        FROM   address
                        WHERE  a.pin_code = ’560012’);
R4: Remove unnecessary joins between a primary key and a foreign key.
Schema
CREATE TABLE products
     p_product_id INTEGER NOT NULL,
     PRIMARY KEY (p_product_id)
CREATE TABLE fact_sales
     f_sales_id   INTEGER NOT NULL,
     f_units_sold INTEGER NOT NULL,
     f_product_id INTEGER NOT NULL,
     PRIMARY KEY (f_sales_id),
     FOREIGN KEY (f_product_id) REFERENCES products(p_product_id)
Original Query
SELECT p_product_id,
       f_units_sold
FROM   fact_sales,
       products
WHERE  f_product_id = p_product_id;
Rewritten Query
SELECT f_product_id,
       f_units_sold
FROM   fact_sales;
R5: Choose EXIST or IN based on subquery selectivity.
Original Query
Select item.id, item.code, item.price
from item
where item.sourceid in (
    Select element.sourceid
    from element
    where element.zip > 1100
order by item.id;
Statistics
Selectivity of different predicates is given below :
( 1 ) source_id > 1100 on table element :: 0.7385
Rewritten Query
Select item.id, item.code, item.price
from item
where exists(select 1
    from element
    where item.sourceid = element.sourceid
      and element.sourceid > 1100
order by item.id;
R6: Pre-filter tables that are involved in self-joins and have low selectivities on their filter and/or join predicates. Remove any redundant filters from the main query. Do not create explicit join statements.
Original Query
with total_price_cte as (
    select item.id, colour.colorcode, sum(item.price) total_price
    from item, color
    where item.colorcode = colour.colorcode
    group by item.id, colour.colorcode
select t_sec.id, t_first.colorcode
from total_price_cte t_first, total_price_cte t_sec
where t_sec.id = t_first.id
    and t_first.colorcode = ‘R’
    and t_sec.colorcode = ‘R’
    and t_first.total_price > 0
order by t_sec.id
limit 100;
Statistics
Selectivity of different predicates is given below :
( 1 ) colour.colorcode = ‘R’ :: 0.01
Rewritten Query
with total_price_cte as (
    select item.id, colour.colorcode, sum(item.price) total_price
    from item, color
    where item.colorcode = colour.colorcode
    group by item.id, colour.colorcode
filtered_total_price_cte as (
    select * from total_price_cte
    where colorcode = ‘R’
select t_sec.id, t_first.colorcode
from filtered_total_price_cte t_first, filtered_total_price_cte t_sec
where t_sec.id = t_first.id
    and t_first.total_price > 0
order by t_sec.id
limit 100;